OpenStudy (anonymous):

lim x→0 cos 9x − cos^2 *9x / 9x

OpenStudy (anonymous):

\[lim~x\rightarrow \infty\{ \cos(9x) -\frac{\cos^2(9x)}{9x}~\text{?}\]

OpenStudy (anonymous):

on the homework its setup like lim x -> 0 ........ cos(9x) - cos^2 9x / 9x

OpenStudy (campbell_st):

break it into smaller parts... 1. \[\lim_{x \rightarrow \infty} \cos(9x) - \lim_{x \rightarrow \infty} \cos^2(9x)/9x\]

OpenStudy (anonymous):

okay, well you can separate this into two distinct limits, one is very easy to deal with because we know what it turns out to be: like this \[lim~ x \rightarrow 0~~cos(9x) - lim~x \rightarrow 0~~\frac{cos^2(9x)}{9x)}\] the other one resembles the sin function you asked about a few questions ago, the same methods should work

OpenStudy (anonymous):

atleast I think it was you that asked about it

OpenStudy (anonymous):

so should it be 1/9 0r 1/81

OpenStudy (campbell_st):

the cos^2(9x) may use l'hopitals rule...

OpenStudy (anonymous):

im thinking the 9's match like cos(9x)/ 9x = 1 and then cos 9x *cos 9x / 9x = 1 am i on the right track?

OpenStudy (anonymous):

campbell, no you can't you don't have a case of 0/0,0/infinity,infinity/0 or infinity/infinity, you have 1/0

OpenStudy (campbell_st):

well lim of cos(9x) = 1 as x ==> 0

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

if you try plotting this one you can see pretty clearly that as x -> 0 the function approaches infinity. there's another pretty useful property of limits that says the limit of a product of two functions is the product of the limits. so what you'd have in this case is \[1-\{lim~x\rightarrow 0~(\cos^2(x))\}*\{lim~x\rightarrow0(\frac{1}{9x})\}\] the limit of one of these is one, while the other is infinity, and 1 times infinity is infinity.

OpenStudy (anonymous):

I dont think its infinity, i put it into the thing at it said an x, but i think its 3

OpenStudy (anonymous):

but its 9(x)

OpenStudy (campbell_st):

use the idea that that the sum of the limits \[\cos ^{2}(9x) = 1 - \sin^2(9x)\] then is \[\lim_{x \rightarrow \infty} \cos(9x) - \lim_{x \rightarrow \infty} (1 - \sin^2(9x))/9x\]

OpenStudy (campbell_st):

\[\lim_{x \rightarrow \infty} \cos(9x) - \lim_{x \rightarrow \infty} 1/9x + \lim_{x \rightarrow \infty} \sin^2(9x)/9x\]